OpenStudy (anonymous):

Explain, in complete sentences, the effect of the difference of squares pattern on the multiplication of radicals. Give an example.

OpenStudy (campbell_st):

it will eliminate the readical and leave an integer e.g. \[(\sqrt{3} + 4)(\sqrt{3} - 4) = (\sqrt{3})^2 - 16 .... = 3 - 16 ... = -13\]